You bought one of Mastadon Manufacturing Co.’s 8 percent coupon bonds one year ago for $1,059. These bonds make annual payments, mature ten years from now, and have a par value of $1,000. Suppose you decide to sell your bonds today, when the required return on the bonds is 7 percent. If the inflation rate was 4.8 percent over the past year, what would be your total real return on the investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Total real return_____%
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text General ta copy. в 1 프 .-. a-Δ. : r_一 逻锂函Merge & Center. $ % , to: conditional Format eCell Insert Delete Format Paste Sort &Find & 2 ClearFe Select Editing Format Painter Formatting, as Table w styles. Styles ▼ ㆆ ▼ Clipboard Font Alignment Number Cells EN28 EJ EK EL EM EN EO EP 21 COUPON LIFE FV YTM PV TODAY 8.00% 10 YEARS 23 24 25 26 27 28 29 30 31 32 1000 7% $1,070.24 FORMULA-PV(796, 10,-80,-1000) PURCHASE PRICE- $1,059.00 TOTAL RETURN (P1+ C1 -PO)/Po 100 TOTAL RETURN = (1070.24 + 80-1059)/1059 100 TOTAL RETURN 8.6157% INFLATION RATE 4.80% 34 35 36 37 38 39 (1+ NOMINAL RATE) (1+ REAL RATE) *(1 INFLATION RATE) (1+0.086157) (1REAL RATE) (1 0.048) 1 REAL RATE 1.086157/1.0481.036409351 REAL RATE = 0.036409351 = 3.64% YIELD SPOT Sheet2 AFN BANKING NOTE BIDDING, UNDERWRITING EUAC AW PM LIFE LP I MRP. INFLATION YIELD WARRANT fund bond CLEAN INVOICE PRICE LI 0:05 31-01-2019